Consider a population consisting of 2,5,6,8, and 10. Suppose samples of size 2 are drawn from this population.



Q1.Compute the mean and the standard deviation of the population.


Q2. List all samples of size 5 and compute the mean for each sample.


Q3. Construct the sampling distribution of the sample means.


Q4. Calculate the mean of the sampling distribution of the sample means.


Q5. Calculate the standard deviation of the sampling distribution of the sample means.

1
Expert's answer
2022-04-29T05:27:54-0400

Q1. We have population values 2,5,6,8,10, population size N=5 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{2+5+6+8+10}{5}=6.2"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}"


"=\dfrac{17.64+1.44+0.04+3.24+14.44}{5}=7.36"

"\sigma=\sqrt{\sigma^2}=\sqrt{7.36}\approx2.712932"

Q2. The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_3=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 2,5,6 & 13/3 \\
 \hdashline
 2 & 2,5,8 & 15/3 \\
 \hdashline
 3 & 2,5,10 & 17/3 \\
 \hdashline
 4 & 2,6,8 & 16/3 \\
 \hdashline
 5 & 2,6,10 & 18/3 \\
 \hdashline
 6 & 2,8,10 & 20/3 \\
 \hdashline
 7 & 5,6,8 & 19/3 \\
 \hdashline
 8 & 5,6,10 & 21/3 \\
 \hdashline
 9 & 5,8,10 & 23/3 \\
 \hdashline
 10 & 6,8,10 & 24/3 \\
 \hdashline 
\end{array}"


Q3.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 13/3 & 1/10 & 13/30 & 169/90 \\
 \hdashline
 15/3 & 1/10 & 15/30 & 225/90 \\
 \hdashline
 16/3 & 1/10 & 16/30 & 256/90 \\
 \hdashline
 17/3 & 1/10 & 17/30 & 289/90 \\
 \hdashline
 18/3 & 1/10 & 18/30 & 324/90 \\
 \hdashline
 19/3 & 1/10 & 19/30 & 361/90 \\
 \hdashline
 20/3 & 1/10 & 20/30 & 400/90 \\
 \hdashline
 21/3 & 1/10 & 21/30 & 441/90 \\
 \hdashline
 23/3 & 1/10 & 23/30 & 529/90 \\
 \hdashline
 24/3 & 1/10 & 24/30 & 576/90 \\
 \hdashline
\end{array}"



Q4. Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=6.2=\mu"



Q5.The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{119}{3}-(6.2)^2=\dfrac{3.68}{3}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{\dfrac{3.68}{3}}\approx1.107550"